# A Quasi-local, Functional Analytic Detection Method for Stationary Limit Surfaces of Black Hole Spacetimes

### v2

@@ III.1 Derivation
To locate the stationary limit surfaces of black hole spacetimes with stationary regions through ellipticity-hyperbolicity transitions of the Hamiltonians of the elementary fermions and bosons, the behaviors of the determinantial polynomials of the principal symbols of these operators have to be analyzed. Thus, we begin by computing the principal symbol (see Equation (1) in Definition II.2) of the Dirac Hamiltonian $\mathcal{H}_{\mathcal{D}}$ defined in Equation (6) at a point $\boldsymbol{y}\in\mathfrak{N}$ of a spacelike hypersurface of a black hole spacetime with stationary regions, which yields
[…]
@@
$(\gamma^{t})^{-1}\,(\gamma^{t})^{-1}=\frac{\mbox{\rm 1 1}_{S_{\boldsymbol{x}}\mathfrak{M}}}{g^{tt}}\quad\textnormal{and}\quad\gamma^{j}\xi_{j}\,\gamma^{k}\xi_{k}=g^{jk}\xi_{j}\xi_{k}\,\mbox{\rm 1 1}_{S_{\boldsymbol{x}}\mathfrak{M}}\,,$
@@
$\bigl{(}\sigma_{(\mathcal{H}_{\textnormal{KG}})_{jk}}(\boldsymbol{y},\boldsymbol{\xi})\bigr{)}_{j,k=1}^{2}=\frac{1}{g^{tt}}\begin{pmatrix}0&g^{tt}\\ g^{jk}\xi_{j}\xi_{k}&-2i\hskip 0.85355ptg^{tk}\xi_{k}\end{pmatrix},$
@@
for all $\boldsymbol{\xi}\in T^{\star}_{\boldsymbol{y}}\mathfrak{N}\backslash\{\boldsymbol{0}\}$ (see Inequality (2) in Definition II.3). However, if this is not the case and if for some $\boldsymbol{\eta}\in T^{\star}_{\boldsymbol{y}}\mathfrak{N}\backslash\{\boldsymbol{0}\}$ satisfying Inequality (13) the two zeros of the second-order polynomial in $\lambda$
@@
Consequently, as the Hamiltonians are elliptic in stationary spacetime regions and hyperbolic in nonstationary spacetime regions, the simultaneous violations of Inequalities (13) and (14), which, taken together, are sufficient conditions for the localization of ellipticity-hyperbolicity transitions, can be used to detect the stationary limit surfaces of a given black hole spacetime with stationary regions. [^4] Since the purely spatial part of the inverse spacetime metric $\boldsymbol{g}^{-1}$ is in general indefinite, this can also be seen from the fact that Equation (15) acts as a stationary limit condition on $\mathfrak{N}$., [^5] Because of their generality, Inequalities (13) and (14) are not limited to black hole spacetimes. Accordingly, they can be employed to locate the stationary limit surfaces of any spacetime having the product structure $\mathfrak{M}=\mathbb{R}\times\mathfrak{N}$. This functional analytic detection method for stationary limit surfaces is quasi-local in the sense that it employs solely operators on $\mathfrak{N}$. Therefore, it locates stationary limit surfaces only on the spacelike hypersurfaces of black hole spacetimes where these operators are defined. Moreover, since the method is based on their principal symbols, it is independent of any particular choice of coordinates on $\mathfrak{N}$, or rather on the associated patch of $\mathfrak{M}$ as long as one uses coordinate systems for which the Hamiltonians are time-independent and coordinate transformations that are diffeomorphisms from $\mathfrak{N}$ to itself. Employing coordinate systems that give rise to time-dependent Hamiltonians, such as Kruskal–Szekeres-type coordinates or Lemaître-type coordinates, [^6] For the original Kruskal–Szekeres and Lemaître coordinates see Kruskal ; Szekeres and Lemaitre , respectively, and for certain generalizations thereof (ChandraBook, , Chapter 6.58.) and Sorge . would, however, break the coordinate invariance and may result in different classifications of the Hamiltonians.
@@ III.2 Application
We now locate the stationary limit surfaces of the Kerr–Newman, Schwarzschild–de Sitter, and Taub–NUT black hole spacetimes using our functional analytic detection method. More precisely, for each case study, we determine and evaluate the violations of Inequalities (13) and (14), thus identifying the locations where the Hamiltonians undergo ellipticity-hyperbolicity transitions.
@@
$\begin{split}\boldsymbol{g}&=\biggl{[}1-\frac{2Mr-Q^{2}}{\Sigma}\biggr{]}\textnormal{d}\tau\otimes\textnormal{d}\tau-\frac{2Mr-Q^{2}}{\Sigma}\,\bigl{(}\textnormal{d}\tau\otimes\bigl{[}\textnormal{d}r-a\sin^{2}(\vartheta)\,\textnormal{d}\varphi\bigr{]}+\bigl{[}\textnormal{d}r-a\sin^{2}(\vartheta)\,\textnormal{d}\varphi\bigr{]}\otimes\textnormal{d}\tau\bigr{)}\\[7.11317pt] &\hskip 12.80365pt-\biggl{[}1+\frac{2Mr-Q^{2}}{\Sigma}\biggr{]}\big{(}\textnormal{d}r-a\sin^{2}(\vartheta)\,\textnormal{d}\varphi\bigr{)}\otimes\big{(}\textnormal{d}r-a\sin^{2}(\vartheta)\,\textnormal{d}\varphi\bigr{)}-\Sigma\,\boldsymbol{g}_{S^{2}}\,,\end{split}$ (16)
where $M\in\mathbb{R}_{>0}$ is the (ADM) mass, $a\in\mathbb{R}$ is the angular momentum per unit mass, and $Q\in\mathbb{R}$ is the charge of the Kerr–Newman black hole satisfying $||(a,Q)||_{\mathbb{R}^{2}}<M$. Moreover, $\Sigma=\Sigma(r,\vartheta):=r^{2}+a^{2}\cos^{2}(\vartheta)$ and $\boldsymbol{g}_{S^{2}}=\textnormal{d}\vartheta\otimes\textnormal{d}\vartheta+\sin^{2}(\vartheta)\,\textnormal{d}\varphi\otimes\textnormal{d}\varphi$ is the metric on the unit $2$-sphere. Substituting the corresponding inverse induced metric on $\mathfrak{N}$ given by
@@
into Inequalities (13) and (14), [^7] Being restricted to coordinate systems for which the purely spatial part of the inverse spacetime metric $\boldsymbol{g}^{-1}$ on $\mathfrak{M}$ is time-independent (cf. the final paragraph of Section III.1 together with Equations (11) and (12)), we can, for the evaluation of these inequalities, use the inverse induced metric $\boldsymbol{g}^{-1}_{\mathfrak{N}}=\phi^{\star}\boldsymbol{g}^{-1}$ on the spacelike hypersurface $\mathfrak{N}$, where $\phi:\mathfrak{N}\hookrightarrow\mathfrak{M}$ is the trivial isometric embedding and the asterisk denotes the pullback from $\mathfrak{M}$ to $\mathfrak{N}$, instead. we obtain, after simple algebraic manipulations,
$[r^{2}-2Mr+a^{2}\cos^{2}(\vartheta)+Q^{2}]\,\xi^{2}_{r}+[a\sin(\vartheta)\,\xi_{r}+\textnormal{csc}(\vartheta)\,\xi_{\varphi}]^{2}+\xi^{2}_{\vartheta}\not=0$ (17)
@@
$\begin{split}&-\bigl{[}r^{2}-2Mr+a^{2}\cos^{2}(\vartheta)+Q^{2}\bigr{]}\bigl{[}\sin^{2}(\vartheta)\,(\eta_{r}\xi_{\vartheta}-\eta_{\vartheta}\xi_{r})^{2}+(\eta_{r}\xi_{\varphi}-\eta_{\varphi}\xi_{r})^{2}\bigr{]}\\[7.11317pt] &\geq\bigl{[}\eta_{\vartheta}\xi_{\varphi}-\eta_{\varphi}\xi_{\vartheta}-a\sin^{2}(\vartheta)\,(\eta_{r}\xi_{\vartheta}-\eta_{\vartheta}\xi_{r})\bigr{]}^{2}\,,\end{split}$ (18)
@@
which are the radial locations of the outer and inner ergosurfaces. Hence, the outer and inner ergosurfaces at ×$\mathbb{R}\times\{r^{\pm}_{\mathcal{E}}\}\times(0,\pi)\times[0,2\pi)$ are the stationary limit surfaces of the Kerr–Newman black hole spacetime. Furthermore, the Hamiltonians are elliptic outside the ergosphere in $(0,r^{-}_{\mathcal{E}})\cup(r^{+}_{\mathcal{E}},\infty)\times(0,\pi)\times[0,2\pi)$ and hyperbolic inside the ergosphere in $(r^{-}_{\mathcal{E}},r^{+}_{\mathcal{E}})\times(0,\pi)\times[0,2\pi)$ with respect to any $\boldsymbol{\eta}\in T^{\star}_{\boldsymbol{y}}\mathfrak{N}\backslash\{\boldsymbol{0}\}$ satisfying Inequality (17) and Inequality (18) for all $\boldsymbol{\xi}\in T^{\star}_{\boldsymbol{y}}\mathfrak{N}\backslash\{\boldsymbol{0}\}$. It is important to stress that the ellipticity-hyperbolicity transitions of the Hamiltonians at the outer and inner ergosurfaces are not artifacts of coordinate singularities, as the nonextreme Kerr–Newman metric in the advanced Eddington–Finkelstein coordinate representation (16) is regular on the entire domain of the black hole spacetime.
@@
$-r\biggl{[}r-2M-\frac{\Lambda r^{3}}{3}\biggr{]}\bigl{[}\sin^{2}(\vartheta)\,(\eta_{r}\xi_{\vartheta}-\eta_{\vartheta}\xi_{r})^{2}+(\eta_{r}\xi_{\varphi}-\eta_{\varphi}\xi_{r})^{2}\bigr{]}\geq(\eta_{\vartheta}\xi_{\varphi}-\eta_{\varphi}\xi_{\vartheta})^{2}\,.$ (20)
Thus, the locations of the ellipticity-hyperbolicity transitions of the Hamiltonians, taking into account all $\boldsymbol{\xi},\boldsymbol{\eta}\in T^{\star}_{\boldsymbol{y}}\mathfrak{N}\backslash\{\boldsymbol{0}\}$, are given by the zeros of the depressed cubic polynomial $r^{3}-3r/\Lambda+6M/\Lambda$, where both of these inequalities do not hold. In more detail, for the physically relevant case $9M^{2}\Lambda\in(0,1)$, there exist, on the one hand, two distinct positive, real-valued zeros at
@@
which, however, lies outside the range of the radial coordinate. Accordingly, the two event horizons at ×$\mathbb{R}\times\{r_{\textnormal{EH, CH}}\}\times(0,\pi)\times[0,2\pi)$ are the stationary limit surfaces of the Schwarzschild–de Sitter black hole spacetime. Besides, the Hamiltonians are elliptic in the region $(r_{\textnormal{EH}},r_{\textnormal{CH}})\times(0,\pi)\times[0,2\pi)$ between the horizons and hyperbolic inside the Schwarzschild-type black hole event horizon and outside the de Sitter-type cosmological event horizon in $(0,r_{\textnormal{EH}})\cup(r_{\textnormal{CH}},\infty)\times(0,\pi)\times[0,2\pi)$ with respect to any $\boldsymbol{\eta}\in T^{\star}_{\boldsymbol{y}}\mathfrak{N}\backslash\{\boldsymbol{0}\}$ fulfilling both Inequality (19) and Inequality (20) for all $\boldsymbol{\xi}\in T^{\star}_{\boldsymbol{y}}\mathfrak{N}\backslash\{\boldsymbol{0}\}$.
@@
$\begin{split}\boldsymbol{g}&=4l^{2}\biggl{[}1-\frac{2(Mr+l^{2})}{r^{2}+l^{2}}\biggr{]}\bigl{(}\textnormal{d}\tau-\cos(\vartheta)\,\textnormal{d}\varphi\bigr{)}\otimes\bigl{(}\textnormal{d}\tau-\cos(\vartheta)\,\textnormal{d}\varphi\bigr{)}-2l\bigl{[}\bigl{(}\textnormal{d}\tau-\cos(\vartheta)\,\textnormal{d}\varphi\bigr{)}\otimes\textnormal{d}r\\[7.11317pt] &\hskip 12.80365pt+\textnormal{d}r\otimes\bigl{(}\textnormal{d}\tau-\cos(\vartheta)\,\textnormal{d}\varphi\bigr{)}\bigr{]}-[r^{2}+l^{2}]\,\boldsymbol{g}_{S^{2}}\,,\end{split}$
@@
$[r^{2}-2Mr-l^{2}]\,\xi^{2}_{r}+\xi^{2}_{\vartheta}+\textnormal{csc}^{2}(\vartheta)\,\xi^{2}_{\varphi}\not=0$ (21)
@@
$-[r^{2}-2Mr-l^{2}]\,\bigl{[}\sin^{2}(\vartheta)\,(\eta_{r}\xi_{\vartheta}-\eta_{\vartheta}\xi_{r})^{2}+(\eta_{r}\xi_{\varphi}-\eta_{\varphi}\xi_{r})^{2}\bigr{]}\geq(\eta_{\vartheta}\xi_{\varphi}-\eta_{\varphi}\xi_{\vartheta})^{2}\,.$ (22)
@@
at which they are not satisfied and therefore the Hamiltonians undergo ellipticity-hyperbolicity transitions. Hence, the Taub–NUT black hole spacetime has stationary limit surfaces at ×$\mathbb{R}\times\{r_{l}^{\pm}\}\times(0,\pi)\times[0,2\pi)$. Moreover, in the region $(-\infty,r_{l}^{-})\cup(r_{l}^{+},\infty)\times(0,\pi)\times[0,2\pi)$ the Hamiltonians are elliptic, whereas in the region $(r_{l}^{-},r_{l}^{+})\times(0,\pi)\times[0,2\pi)$ they are hyperbolic with respect to any $\boldsymbol{\eta}\in T^{\star}_{\boldsymbol{y}}\mathfrak{N}\backslash\{\boldsymbol{0}\}$ satisfying Inequality (21) and Inequality (22) for all $\boldsymbol{\xi}\in T^{\star}_{\boldsymbol{y}}\mathfrak{N}\backslash\{\boldsymbol{0}\}$.

### v3

@@ III.1 Derivation
To locate the stationary limit surfaces of black hole spacetimes with stationary regions through ellipticity-hyperbolicity transitions of the Hamiltonians of the elementary fermions and bosons, the behaviors of the determinantial polynomials of the principal symbols of these operators have to be analyzed. Thus, we begin by computing the principal symbol [see Equation (1) in Definition II.2] of the Dirac Hamiltonian $\mathcal{H}_{\mathcal{D}}$ defined in Equation (6) at a point $\boldsymbol{y}\in\mathfrak{N}$ of a spacelike hypersurface of a black hole spacetime with stationary regions, which yields
$\sigma_{\mathcal{H}_{\mathcal{D}}}(\boldsymbol{y},\boldsymbol{\xi})=-i\hskip 1.13791pt(\gamma^{t})^{-1}\gamma^{j}\xi_{j}\,$,
@@
$(\gamma^{t})^{-1}\,(\gamma^{t})^{-1}=\frac{\mbox{\rm 1 \hskip-10.50005pt 1}_{S_{\boldsymbol{x}}\mathfrak{M}}}{g^{tt}}\quad\textnormal{and}\quad\gamma^{j}\xi_{j}\,\gamma^{k}\xi_{k}=g^{jk}\xi_{j}\xi_{k}\,\mbox{\rm 1 \hskip-10.50005pt 1}_{S_{\boldsymbol{x}}\mathfrak{M}}\,$,
@@
$\bigl(\sigma_{(\mathcal{H}_{\textnormal{KG}})_{jk}}(\boldsymbol{y},\boldsymbol{\xi})\bigr)_{j,k=1}^{2}=\frac{1}{g^{tt}}\hskip-1.42271pt\begin{pmatrix}0&g^{tt}\\ g^{jk}\xi_{j}\xi_{k}&-2i\hskip 0.85355ptg^{tk}\xi_{k}\end{pmatrix}$,
@@
for all $\boldsymbol{\xi}\in T^{\star}_{\boldsymbol{y}}\mathfrak{N}\backslash\{\boldsymbol{0}\}$ [see Inequality (2) in Definition II.3]. However, if this is not the case and if for some $\boldsymbol{\eta}\in T^{\star}_{\boldsymbol{y}}\mathfrak{N}\backslash\{\boldsymbol{0}\}$ satisfying Inequality (13) the two zeros of the second-order polynomial in $\lambda$
@@
Consequently, as the Hamiltonians are elliptic in stationary spacetime regions and hyperbolic in nonstationary spacetime regions, the simultaneous violations of Inequalities (13) and (14), which, taken together, are sufficient conditions for the localization of ellipticity-hyperbolicity transitions, can be used to detect the stationary limit surfaces of a given black hole spacetime with stationary regions. [^4] Since the purely spatial part of the inverse spacetime metric $\boldsymbol{g}^{-1}$ is in general indefinite, this can also be seen from the fact that Equation (15) acts as a stationary limit condition on $\mathfrak{N}$., [^5] Because of their generality, Inequalities (13) and (14) are not limited to black hole spacetimes. Accordingly, they can be employed to locate the stationary limit surfaces of any spacetime having the product structure $\mathfrak{M}=\mathbb{R}\times\mathfrak{N}$. This functional analytic detection method for stationary limit surfaces is quasi-local in the sense that it employs solely operators on $\mathfrak{N}$. Therefore, it locates stationary limit surfaces only on the spacelike hypersurfaces of black hole spacetimes where these operators are defined. Moreover, since the method is based on their principal symbols, it is independent of any particular choice of coordinates on $\mathfrak{N}$, or rather on the associated patch of $\mathfrak{M}$ as long as one uses coordinate systems for which the Hamiltonians are time-independent and coordinate transformations that are diffeomorphisms from $\mathfrak{N}$ to itself. Working with coordinate systems that give rise to time-dependent Hamiltonians, such as Kruskal–Szekeres-type coordinates or Lemaître-type coordinates, [^6] For the original Kruskal–Szekeres and Lemaître coordinates see Kruskal ; Szekeres and Lemaitre , respectively, and for certain generalizations thereof (ChandraBook, , Chapter 6.58.) and Sorge . would, however, break the coordinate invariance and may result in different classifications of the Hamiltonians. Hence, being limited to coordinate systems on $\mathfrak{M}$ for which the purely spatial part $(g^{ij})_{i,j\in\{1,2,3\}}$ of the inverse spacetime metric $\boldsymbol{g}^{-1}$ is time-independent, [^7] This directly follows from Equations (11) and (12) together with the time-independence of the Hamiltonians. Inequalities (13) and (14) can, instead of $(g^{ij})_{i,j\in\{1,2,3\}}$, be evaluated by the inverse induced metric $\boldsymbol{g}^{-1}_{\mathfrak{N}}=\phi^{\star}\boldsymbol{g}^{-1}$ on the spacelike hypersurface $\mathfrak{N}$, where $\phi\hskip-1.9919pt:\mathfrak{N}\hookrightarrow\mathfrak{M}$ is the trivial isometric embedding and the asterisk denotes the pullback from $\mathfrak{M}$ to $\mathfrak{N}$. This aspect further illustrates the quasi-locality of the functional analytic detection method.
@@ III.2 Application
We now locate the stationary limit surfaces of the Kerr–Newman, Schwarzschild–de Sitter, and Taub–NUT black hole spacetimes using our functional analytic detection method. More precisely, for each case study, we determine and examine the violations of Inequalities (13) and (14), thus identifying the locations where the Hamiltonians undergo ellipticity-hyperbolicity transitions.
@@
[…]
where $M\in\mathbb{R}_{>0}$ is the (ADM) mass, $a\in\mathbb{R}$ is the angular momentum per unit mass, and $Q\in\mathbb{R}$ is the charge of the Kerr–Newman black hole satisfying $(a,Q)||_{\mathbb{R}^{2}}<M$. Moreover, $\Sigma=\Sigma(r,\vartheta):=r^{2}+a^{2}\cos^{2}(\vartheta)$ and $\boldsymbol{g}_{S^{2}}=\textnormal{d}\vartheta\otimes\textnormal{d}\vartheta+\sin^{2}(\vartheta)\,\textnormal{d}\varphi\otimes\textnormal{d}\varphi$ is the metric on the unit $2$-sphere. Substituting the corresponding inverse induced metric on $\mathfrak{N}$ given by
@@
into Inequalities (13) and (14), we obtain, after simple algebraic manipulations,
$r^{2}-2Mr+a^{2}\cos^{2}(\vartheta)+Q^{2}]\,\xi^{2}_{r}+[a\sin(\vartheta)\,\xi_{r}+\textnormal{csc}(\vartheta)\,\xi_{\varphi}]^{2}+\xi^{2}_{\vartheta}\not=0$ (17)
@@
[…]
@@
which are the radial locations of the outer and inner ergosurfaces. Hence, the outer and inner ergosurfaces at $\mathbb{R}\times\{r^{\pm}_{\mathcal{E}}\}\times(0,\pi)\times[0,2\pi)$ are the stationary limit surfaces of the Kerr–Newman black hole spacetime. Furthermore, the Hamiltonians are elliptic outside the ergosphere in $(0,r^{-}_{\mathcal{E}})\cup(r^{+}_{\mathcal{E}},\infty)\times(0,\pi)\times[0,2\pi)$ and hyperbolic inside the ergosphere in $(r^{-}_{\mathcal{E}},r^{+}_{\mathcal{E}})\times(0,\pi)\times[0,2\pi)$ with respect to any $\boldsymbol{\eta}\in T^{\star}_{\boldsymbol{y}}\mathfrak{N}\backslash\{\boldsymbol{0}\}$ satisfying Inequality (17) and Inequality (18) for all $\boldsymbol{\xi}\in T^{\star}_{\boldsymbol{y}}\mathfrak{N}\backslash\{\boldsymbol{0}\}$. It is important to stress that the ellipticity-hyperbolicity transitions of the Hamiltonians at the outer and inner ergosurfaces are not artifacts of coordinate singularities, as the nonextreme Kerr–Newman metric in the advanced Eddington–Finkelstein coordinate representation (16) is regular on the entire domain of the black hole spacetime.
@@
$r\biggl[r-2M-\frac{\Lambda r^{3}}{3}\biggr]\bigl[\sin^{2}(\vartheta)\,(\eta_{r}\xi_{\vartheta}-\eta_{\vartheta}\xi_{r})^{2}+(\eta_{r}\xi_{\varphi}-\eta_{\varphi}\xi_{r})^{2}\bigr]\geq(\eta_{\vartheta}\xi_{\varphi}-\eta_{\varphi}\xi_{\vartheta})^{2}\,.$ (20)
Thus, the locations of the ellipticity-hyperbolicity transitions of the Hamiltonians, taking into account all $\boldsymbol{\xi},\boldsymbol{\eta}\in T^{\star}_{\boldsymbol{y}}\mathfrak{N}\backslash\{\boldsymbol{0}\}$, are given by the zeros of the depressed cubic polynomial $r^{3}-3r/\Lambda+6M/\Lambda$ where both of these inequalities do not hold. In more detail, for the physically relevant case $9M^{2}\Lambda\in(0,1)$, there exist, on the one hand, two distinct positive, real-valued zeros at
@@
which, however, lies outside the range of the radial coordinate. Accordingly, the two event horizons at $\mathbb{R}\times\{r_{\textnormal{EH, CH}}\}\times(0,\pi)\times[0,2\pi)$ are the stationary limit surfaces of the Schwarzschild–de Sitter black hole spacetime. Besides, the Hamiltonians are elliptic in the region $(r_{\textnormal{EH}},r_{\textnormal{CH}})\times(0,\pi)\times[0,2\pi)$ between the horizons and hyperbolic inside the Schwarzschild-type black hole event horizon and outside the de Sitter-type cosmological event horizon in $(0,r_{\textnormal{EH}})\cup(r_{\textnormal{CH}},\infty)\times(0,\pi)\times[0,2\pi)$ with respect to any $\boldsymbol{\eta}\in T^{\star}_{\boldsymbol{y}}\mathfrak{N}\backslash\{\boldsymbol{0}\}$ fulfilling both Inequality (19) and Inequality (20) for all $\boldsymbol{\xi}\in T^{\star}_{\boldsymbol{y}}\mathfrak{N}\backslash\{\boldsymbol{0}\}$.
@@
[…]
@@
$r^{2}-2Mr-l^{2}]\,\xi^{2}_{r}+\xi^{2}_{\vartheta}+\textnormal{csc}^{2}(\vartheta)\,\xi^{2}_{\varphi}\not=0$ (21)
@@
$r^{2}-2Mr-l^{2}]\,\bigl[\sin^{2}(\vartheta)\,(\eta_{r}\xi_{\vartheta}-\eta_{\vartheta}\xi_{r})^{2}+(\eta_{r}\xi_{\varphi}-\eta_{\varphi}\xi_{r})^{2}\bigr]\geq(\eta_{\vartheta}\xi_{\varphi}-\eta_{\varphi}\xi_{\vartheta})^{2}\,.$ (22)
@@
at which they are not satisfied and therefore the Hamiltonians undergo ellipticity-hyperbolicity transitions. Hence, the Taub–NUT black hole spacetime has stationary limit surfaces at $\mathbb{R}\times\{r_{l}^{\pm}\}\times(0,\pi)\times[0,2\pi)$. Moreover, in the region $(-\infty,r_{l}^{-})\cup(r_{l}^{+},\infty)\times(0,\pi)\times[0,2\pi)$ the Hamiltonians are elliptic, whereas in the region $(r_{l}^{-},r_{l}^{+})\times(0,\pi)\times[0,2\pi)$ they are hyperbolic with respect to any $\boldsymbol{\eta}\in T^{\star}_{\boldsymbol{y}}\mathfrak{N}\backslash\{\boldsymbol{0}\}$ satisfying Inequality (21) and Inequality (22) for all $\boldsymbol{\xi}\in T^{\star}_{\boldsymbol{y}}\mathfrak{N}\backslash\{\boldsymbol{0}\}$.